wo pucks collide on an air hockey table. Puck A has a mass of 26.0 g and is initially traveling in the +xdirection at 6.50 m/s. Puck B has a mass of 78.0 g and is initially at rest. After the pucks collide, puck A moves away at an angle of 58.0$^\circ$ above the +x axis, while puck B travels at an angle of 26.0$^\circ$ below the +x axis.
Calculate puck A's final speed.
Calculate puck B's final speed.
What fraction of kinetic energy was lost in this collision?
Answer #1

Here ,

mass of A , mA = 26 gm

uA = 6.5 i m/s

mass of B . mB = 78 gm

uB = 0 ,

thetaA = 58 degree

thetaB = -26 degree

let the final speed of A is vA

final speed of B is vB

Using conservation of momentum

vA * sin(58) * 26 - vB * 78 * sin(26) = 0

also ,

vA * cos(58) * 26 + vB * cos(26) * 78 = 6.5 * 26

vA = 2.86 m/s

vB = 1.85 m/s

the final speed of A is 2.86 m/s

the final speed of B is 1.85 m/s


fraction of energy lost = (0.5 * 26 * 6.5^2 - 0.5 * 26 * 2.86^2 - 0.5 * 78 * 1.85^2)/(0.5 * 26 * 6.5^2 )

fraction of energy lost = 0.563

the fraction of energy lost is 0.564